Howe Hinges Co manufactures and sells a single product. This product has the following operational data:

Unit sales price $30

Variable cost per unit 18

Fixed costs 111,000

Income tax rate 30%

At a current level of $300,000 in sales, by what percent could sales fall before the company starts losing money?
Answer #1

At 7.5% company sales could fall , before the company starts loosing money.

Explantion and Calculation

Breakeven point is the point below which company will make loss . Hence upto Breakeven point Sales could fall but below that company will start making losses, and this level is known as Margin of Safety

Margin of safety is the point upto which sales can fall before company can start making loss . From Sale point to breakeven point level is called Margin of Safety level. At breakeven point compnay will be at no Profit and loss below breakeven point company will make losses as it will not be able to recover their fixed expenses.

Formula for Margin of Safety (%) = (Current Sales - Breakeven Sales )/ Current Sales × 100

One more way is  

MOS (%) = 100 - BEP Sales (%)

Formula for Break even (Unit )= Fixed Cost / Contribution margin per unit

Contribution Margin per unit = SElling Price - Variable cost per unit

=30-18 =12 Per unit

Break even point unit =111000/12

=9250

Breakeven point Sales = BEP ( unit ) × Selling price per unit

=9250×30

=277500

Margin of Safety percentage = (Sales - BEP Sales )/ Sales ×100

=[(300000-277500)/300000 ]×100

=7.5 %

Hence Upto 7.5% Sales could fall after that is will start loosing money